OpenStudy (jolquinnzley):

PLEASE HELP QUICKLY!! Prove the Converse of the Pythagorean Theorem using similar triangles. The Converse of the Pythagorean Theorem states that when the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse, the triangle is a right triangle. Be sure to create and name the appropriate geometric figures.

OpenStudy (saadsword):

the pythagorean theorem is a^2 + b^2 = c^2, c is the hypotenuse, and a & b are the legs. so the converse is IF c^2 = a^2 + b^2, then the triangle is a right triangle.

OpenStudy (saadsword):

otherwise if a^2 + b^2 does NOT equal c^2, then the triangle is NOT a right triangle.

OpenStudy (jolquinnzley):

Thank you so much! It makes sense to me for once, I appreciate your help!

OpenStudy (saadsword):

no problem!!